# Personalized Federated Learning with Heat-Kernel Enhanced Tensorized Multi-View Clustering

### v1

@@ Abstract
We present a robust personalized federated learning framework that leverages heat-kernel enhanced tensorized multi-view fuzzy c-means clustering with advanced tensor decomposition techniques. Our approach integrates heat-kernel coefficients adapted from quantum field theory with Tucker decomposition and canonical polyadic decomposition (CANDECOMP/PARAFAC) to transform conventional distance metrics and efficiently represent high-dimensional multi-view structures. The framework employs matriculation and vectorization techniques to facilitate the discovery of hidden structures and multilinear relationships via N-way generalized tensors. The proposed method introduces a dual-level optimization scheme: local heat-kernel enhanced fuzzy clustering with tensor decomposition operating on order-N input tensors, and federated aggregation of tensor factors with privacy-preserving personalization mechanisms. The local stage employs tensorized kernel Euclidean distance transformations and Tucker decomposition to discover client-specific patterns in multi-view tensor data, while the global aggregation process coordinates tensor factors (core tensors and factor matrices) across clients through differential privacy-preserving protocols. This tensorized approach enables efficient handling of high-dimensional multi-view data with significant communication savings through low-rank tensor approximations.
Keywords: Federated learning, Multi-view clustering, Heat-kernel methods, Tensor decomposition, PARAFAC2, Tucker decomposition, Personalized machine learning, Privacy-preserving clustering.
@@ 1 Introduction
The emergence of distributed multi-view data across heterogeneous clients presents unique challenges for federated learning systems. Traditional federated clustering assumes homogeneous data representations, failing to address the complexity of multi-view tensor structures where clients may have different modalities, feature spaces, and local data characteristics.
The massive growing of data is happening as the advance and revolution of technology in many fields becomes so forefront and powerful. Machine learning algorithms also developed and extending to significantly enables these data trends improved analysis. Mathematically, a data is representing by a unit or collection of row and column numbers arranged in a rectangular or array-like fashion. An array representing vectors or matrices storage set of different numbers we wish to have. In unsupervised clustering, we partition the rows unit data based on its given dimension attribute into some groups or clusters. This is a typical data that used as an input matrix to reveal important information in an efficient manner. However, a multi-resources data requires its representation to better process using tensor space aiming to meet the significant raise of data growth. Reconstruction tensor-based approaches to represent heterogeneous data can be challenging as it involves multilinear relationship between sets of matrix algebra but promising satisfaction output.
Most common application of third-order tensor representation is a color image. RGB image is stored in a 3-Dimensional array as $\text{channels}\times\text{height}\times\text{width}$. Fig. 1 displayed the illustration of R, G, and B component of images. As it can be seen, the shape of the image in Fig. 1 is $7\times 7\times 3$. The colour channels stored the height and width of the single image color of R, G, and B. Usually, the pixel values of a single image will be arranged in a 1-Dimensional vector to get the required format for classification or clustering process.
@@
Tensor-based approaches had been extended in many machines learning techniques, including K-means clustering. Singular value decomposition (SVD) and principal component analysis (PCA) [1, 2] are two feature reduction techniques which frequently used to construct the data before starting classification or clustering process. Both SVD and PCA can process a data matrix which is well known as tensor of order two. On the other hand, multi-view (MV) data collects information from multiple sources. Each sources represent their feature component about one observation or object in a data. Mathematically, each source can be regarded as view. Accordingly, view represented by matrix space. For instance, each matrix on MV data may represent video recording, images, text, or audio recording’ feature component. A representation of these data sources possibly achieves good performances with tensor-based reconstruction as it can be reshaping based on their multicollinearity. In tensor space, MV data considered as a third-order tensor, with size $n\times d_{h}\times s$.
From the perspective of MV clustering, each component of feature-view stored as a two-order tensor. In this sense, if one MV data contains four views, then MV clustering algorithm will cluster the data by processing these four-two-Dimensional matrix based on their similarities. Most commonly, the similarities are measured by estimating the Euclidean distance between data points and cluster centers within/across views. The analogy of diversity and heterogeneity on multiple representations of the same information is precise [3]. For instance, online media such as CNN and FOX featuring a news about technology or culture based on different observation depending on their own perspective-based, context-based, purpose-based, etc. The different writing styles, image or video, and short description provided is unique (“heterogeneous”) and has semantic relationship. Hence, the representation to determine the space of variations to occur the unity might be useful if the structuring processes involves tensor-based reconstruction forming more smaller matrices and reduce the time complexity of conventional MV clustering algorithms.
The conventional k-means clustering not effective to directly explore structure of tensor space as it is required lots of memory. To fit into memory required processes, we will enforce the amount of data greatly using CANDECOMP (PARAFAC2). Our main motivation in using tensor-based Tucker decomposition approaches is to reveal hidden information that often happening in one data. The elaboration of these decomposition of PARAFAC2 and Tucker model will be also implemented into MV data directing its tensor transformation and modification of k-means objective function for processing higher-order tensor data.
2 Methodology
The integration of federated learning with multi-view clustering and tensor decomposition represents a convergence of several active research domains. This section reviews the foundational work and recent advances in federated learning, multi-view clustering, tensor decomposition methods, and their intersections, positioning our contributions within the broader landscape of distributed machine learning.
@@ 2.1 Federated Learning and Clustering
Federated learning, pioneered by McMahan et al. [4], enables collaborative machine learning across distributed clients without centralizing raw data. The seminal FedAvg algorithm established the foundation for aggregating local model updates through weighted averaging. However, traditional federated learning primarily focuses on supervised tasks, with federated clustering receiving comparatively less attention despite its practical importance in distributed data analytics.
Konečnỳ et al. [5] introduced federated optimization as a method to train a centralized model using decentralized data stored across numerous devices, emphasizing communication efficiency due to limited connectivity and privacy concerns. They proposed modifications to existing algorithms to adapt them to federated optimization settings. Li et al. [6] adopted the foundational ideas of federated optimization and extended them by integrating advanced techniques like gradient compression, acceleration, and variance reduction to address communication bottlenecks and improve the scalability of federated learning systems
Early work on federated clustering by Dennis et al. [7], Ghosh et al. [8], and Makhija et al. [9] demonstrated the feasibility of distributed k-means and hierarchical clustering. These approaches typically involve local clustering followed by global aggregation of cluster centers or assignments. However, these methods assume homogeneous data distributions and fail to address the inherent heterogeneity in federated environments.
Recent advances have introduced personalized federated clustering to address client heterogeneity. Caldarola et al. [10] proposed adaptive weight schemes for heterogeneous clustering, while Li et al. [11] developed meta-learning approaches for personalized cluster initialization. These works, however, are limited to single-view data and do not leverage the rich multi-modal information available in modern federated systems.
Privacy-preserving aspects of federated clustering have been addressed through differential privacy mechanisms [12] and secure aggregation protocols [13]. Geyer et al. [14] demonstrated how differential privacy can be integrated into federated clustering while maintaining acceptable utility. However, these privacy mechanisms often come at the cost of clustering accuracy, particularly in high-dimensional settings.
The communication efficiency challenge in federated clustering has been tackled through gradient compression [15], sparsification techniques [16], and model quantization [17]. While these methods reduce communication overhead, they typically do not exploit the inherent structure of multi-view tensor data for more efficient representation and transmission.
@@ 2.2 Multi-View Clustering
Multi-view clustering leverages complementary information from multiple data representations to improve clustering performance. Early work by Bickel and Scheffer [18] established the theoretical foundations for multi-view learning, demonstrating superior performance when views provide complementary information about the underlying data structure.
@@
Classical multi-view clustering approaches can be categorized into three main paradigms: co-training based methods, subspace learning approaches, and consensus clustering techniques. Co-training methods, such as the work by Kumar et al. [19], iteratively refine clustering results by leveraging inter-view consistency. Subspace learning approaches, including Canonical Correlation Analysis (CCA) based methods [20] and shared subspace clustering [21], project multiple views into a common representation space for unified clustering.
@@
Graph-based multi-view clustering has gained significant attention due to its ability to capture complex inter-view relationships. Wang et al. [22] proposed multi-view graph clustering with adaptive neighbor learning, while Nie et al. [23] developed parameter-free auto-weighted multiple graph learning. These methods construct view-specific graphs and learn optimal view weights for clustering.
@@
Deep learning approaches have revolutionized multi-view clustering through end-to-end representation learning. Zhang et al. [24] introduced auto-encoder based multi-view clustering, while Zhu et al. [25] developed deep multi-view clustering networks with view correlation discovery. However, these methods typically require centralized data access and are not designed for federated environments.
@@
Fuzzy clustering extensions to multi-view scenarios have been explored by Jiang et al. [26] and Wang et al. [27], introducing weighted aggregation schemes and view-specific fuzzification parameters. These soft clustering approaches provide uncertainty quantification but lack the scalability and privacy considerations required for federated deployment.
@@
Our goal is to provide a comprehensive mathematical objective function to robustly handle multi-view data distributed on multiple clients using tensor algebra, with non-linearity compatibilities. Thus, we replace the distance between data points and cluster centers across data views in Euclidean space with a kernel Euclidean distance (KED). We define the mathematical expression of KED in the following way:
@@
In Eq. 5, we treat all data views equally important, which is not efficient to solve multiple resources data. Since different views represents different feature behavior, it is necessary to assign a different weight $v_{h}$ to each data view $x_{i}^{h}$, with a normalization approach $\sum_{h=1}^{s}v_{h}=1$. By considering weight factors $v_{h}$, we can modify Eq. 5 in the following way.

### v2

# Personalized Enhanced Federated Multi-View Clustering via Heat-Kernel Tensor Decomposition

@@ Abstract
This paper proposes a personalized federated learning framework integrating heat-kernel enhanced tensorized multi-view fuzzy c-means clustering with tensor decomposition techniques. The approach combines heat-kernel coefficients adapted from quantum field theory with PARAFAC2 and Tucker decomposition to transform distance metrics and efficiently represent high-dimensional multi-view structures. Two main algorithms, FedHK-PARAFAC2 and FedHK-Tucker, are developed to extract shared and view-specific features while preserving inter-view relationships. The framework addresses data heterogeneity, privacy preservation, and communication efficiency challenges in federated learning environments. Theoretical analysis provides convergence guarantees, privacy bounds, and complexity analysis. The integration of heat-kernel methods with tensor decomposition in a federated setting offers a novel approach for effective multi-view data analysis while ensuring data privacy.
Keywords: Federated learning, Multi-view clustering, Heat-kernel methods, Tensor decomposition, PARAFAC2, Tucker decomposition, Privacy preservation.
@@ 1 Introduction
Tensor-based machine learning approaches have emerged as powerful tools for analyzing multi-dimensional data structures, enabling the capture of complex relationships and interactions across multiple modes. Tensors, as generalizations of matrices to higher dimensions, provide a natural framework for representing multi-view data, where each view corresponds to a different modality or feature set. For instance, in image processing, a color image can be represented as a third-order tensor with dimensions corresponding to height, width, and color channels (Red, Green, Blue) as illustrated in Figure 1.
@@
As can be seen in Figure 1, a color image is represented as a 3-dimensional tensor, where each slice along the third dimension corresponds to one of the RGB color channels. This tensor representation allows for the preservation of spatial and color information, enabling more effective analysis and processing of image data. RGB image is stored in a 3-Dimensional array as $\text{channels}\times\text{height}\times\text{width}$ or $\text{height}\times\text{width}\times\text{channels}$. The $7\times 7\times 3$ tensor in Figure 1 represents a small RGB image with a height and width of 7 pixels and 3 color channels. Each element in the tensor corresponds to the intensity value of a specific color channel at a particular pixel location.
In federated learning environments, clients often possess multi-view (MV) data that can be naturally represented as tensors. For example, in healthcare applications, different hospitals may collect various types of patient data, such as medical images, electronic health records, and genomic data, each representing a different view of the patient’s health status. By leveraging tensor-based representations, federated learning algorithms can effectively capture the multi-modal nature of the data while preserving privacy and minimizing communication overhead. Singular value decomposition (SVD) and principal component analysis (PCA) [welke2006fast, sinaga2021entropy] are classical matrix factorization techniques that have been widely used for dimensionality reduction and feature extraction. However, these methods are limited to two-dimensional data structures and cannot fully exploit the multi-way relationships present in tensor data. Tensor decomposition methods, such as CANDECOMP/PARAFAC (CP) [harshman1970foundations] and Tucker decomposition [tucker1966some], extend these concepts to higher-order tensors, enabling more comprehensive analysis of multi-view data.
From the perspective of MV clustering, each component of feature-view stored as a two-order tensor (matrix) can be naturally represented as a three-order tensor by stacking the matrices along a new dimension corresponding to the views. This tensor representation allows for the preservation of inter-view relationships and facilitates the application of tensor decomposition techniques to extract shared and view-specific features. By leveraging tensor algebra, MV clustering algorithms can effectively capture the complex interactions between different views, leading to improved clustering performance and robustness.
The analogy of diversity and heterogeneity on multiple representations of the same information is precise [giunchiglia2022representation] . In this sense, multi-view data can be viewed as different perspectives or modalities of the same underlying phenomenon, each providing unique insights and complementary information. For example, in social network analysis, user profiles may include text data from posts, images shared, and interaction patterns with other users. Each of these views captures different aspects of user behavior and preferences, contributing to a more holistic understanding of the social dynamics.
This paper presents two novel algorithms, namely FedHK-PARAFAC2 and FedHK-Tucker, that integrate heat-kernel enhanced tensorized multi-view fuzzy c-means clustering with advanced tensor decomposition techniques in a federated learning framework. The proposed methods leverage heat-kernel coefficients adapted from quantum field theory to transform conventional distance metrics, enabling the capture of local geometric relationships within multi-view data distributed across heterogeneous clients. By incorporating PARAFAC2 and Tucker decomposition, the algorithms efficiently represent high-dimensional multi-view structures while addressing challenges of data heterogeneity, privacy preservation, and communication efficiency inherent in federated learning environments.
2 Related Work
In order to provide a comprehensive overview of the contributions of this study, a review of the extant literature was conducted across four key domains: federated learning and clustering, multi-view clustering, heat kernel methods and quantum-inspired learning, and tensor decomposition techniques. This review underscores foundational works and recent advancements in the field, while concurrently identifying gaps that the proposed framework aims to address.
@@ 2.1 Federated Learning and Clustering
Pattern recognition and clustering in federated learning (FL) environments have garnered significant attention due to the increasing need for privacy-preserving distributed data analysis. Federated learning, introduced by McMahan et al. [mcmahan2017communication], enables collaborative model training across decentralized clients without sharing raw data. While initial research focused on supervised learning tasks, recent efforts have extended federated learning to unsupervised clustering scenarios. Early work by Dennis et al. [dennis2021heterogeneity] explored federated k-means clustering, demonstrating the feasibility of distributed clustering with local updates and global aggregation. Ghosh et al. [ghosh2002consensus] proposed consensus-based approaches for federated hierarchical clustering, addressing challenges of data heterogeneity and communication efficiency. Makhija et al. [makhija2022federated] introduced federated fuzzy c-means clustering, leveraging soft assignments to enhance robustness against noise and outliers. However, these methods often assume homogeneous data distributions and do not fully exploit multi-view data structures prevalent in modern applications. Recent advancements have focused on personalized federated clustering to accommodate client heterogeneity. Caldarola et al. [caldarola2021cluster] proposed adaptive weight schemes for heterogeneous clustering, while Li et al. [li2022hpfl] developed meta-learning approaches for personalized cluster initialization. Privacy-preserving aspects have been addressed through differential privacy mechanisms [kairouz2021advances] and secure aggregation protocols [bonawitz2017practical], ensuring that sensitive information remains protected during the clustering process. Communication efficiency has been tackled through gradient compression [konecny2016federated], sparsification techniques [aji2017sparse], and model quantization [reisizadeh2020fedpaq], reducing the overhead associated with transmitting model updates across clients.
@@ 2.2 Multi-View Clustering
Multi-view clustering (MVC) has emerged as a powerful paradigm for leveraging complementary information from multiple data representations to enhance clustering performance. Early work by Bickel and Scheffer [bickel2004multi] established the theoretical foundations for multi-view learning, demonstrating that integrating multiple views can lead to improved clustering accuracy. Classical MVC approaches can be categorized into co-training based methods, subspace learning approaches, and consensus clustering techniques. Co-training methods, such as those proposed by Kumar et al. [kumar2011co], iteratively refine clustering results by leveraging inter-view consistency. Subspace learning approaches, including Canonical Correlation Analysis (CCA) based methods [hotelling1992relations] and shared subspace clustering [wang2019multiple], project multiple views into a common representation space for unified clustering. Graph-based MVC has gained significant attention due to its ability to capture complex inter-view relationships. Wang et al. [wang2019gmc] proposed multi-view graph clustering with adaptive neighbor learning, while Nie et al. [nie2016parameter] developed parameter-free auto-weighted multiple graph learning. Deep learning approaches have revolutionized MVC through end-to-end representation learning. Zhang et al. [zhang2019ae2] introduced auto-encoder based multi-view clustering, while Zhu et al. [zhu2019multi] developed deep multi-view clustering networks with view correlation discovery. Fuzzy clustering extensions to multi-view scenarios have been explored by Jiang et al. [jiang2014collaborative] and Wang et al. [wang2017multi], introducing weighted aggregation schemes and view-specific fuzzification parameters. Sinaga et al. [sinaga2025parameter] proposed a parameter-free multi-view fuzzy clustering algorithm that automatically learns view weights and fuzzification parameters, enhancing clustering robustness and interpretability. However, these methods typically require centralized data access and are not designed for federated environments.
The federated multi-view clustering is first introduced by Chen et al. [chen2022federated] for federated multi-view representation learning and Liu et al. [liu2023federated] for privacy-preserving multi-view clustering. However, these methods do not leverage tensor decomposition for efficient representation and communication. Yang et al. [yang2024federated] proposed a federated multi-view clustering framework that combines local clustering with global consensus, but it lacks the integration of heat-kernel methods and advanced tensor decomposition techniques. Earlier, Sinaga [sinaga2025fedhk] proposed a heat-kernel enhanced federated multi-view clustering framework, but it did not incorporate tensor decomposition methods for efficient multi-view representation. Thus, there is a need for advanced federated multi-view clustering approaches that leverage tensor decomposition and heat-kernel methods to address data heterogeneity, privacy preservation, and communication efficiency challenges in distributed settings.
2.3 Heat Kernel Methods and Quantum-Inspired Learning
Heat kernel methods have emerged as powerful tools for capturing local geometric structure in high-dimensional data spaces. Originally developed in differential geometry and quantum field theory, heat kernels provide a natural way to encode multi-scale geometric information through diffusion processes on manifolds. The heat kernel, defined as the fundamental solution to the heat equation, has found extensive applications in machine learning through kernel methods and manifold learning. Ham et al. [ham2004kernel] introduced diffusion maps that leverage heat kernel properties for dimensionality reduction, while Coifman and Lafon [coifman2006diffusion] established theoretical foundations for diffusion-based data analysis. In clustering applications, heat kernel methods have been used to capture local neighborhood structure and improve cluster boundary detection. Zelnik-Manor and Perona [zelnik2004self] developed self-tuning spectral clustering using local heat kernel scales, while von Luxburg [von2007tutorial] provided comprehensive analysis of heat kernel properties in graph-based clustering.
Quantum-inspired machine learning has gained attention for its ability to leverage quantum mechanical principles in classical algorithms. Quantum kernels, introduced by Havlíček et al. [havlivcek2019supervised], enable the computation of quantum feature maps for enhanced pattern recognition. The connection between quantum field theory and machine learning has been explored through path integral formulations [carleo2017solving] and quantum neural networks [biamonte2017quantum]. The integration of heat kernel coefficients with clustering algorithms represents a natural extension of kernel methods to soft clustering scenarios. By incorporating heat diffusion properties, clustering algorithms can better capture local data manifold structure and improve robustness to noise and outliers.
2.4 Tensor Decomposition Methods
Tensor decomposition has emerged as a powerful paradigm for analyzing high-dimensional multi-modal data, providing compact representations that capture multilinear relationships inherent in complex datasets. The field has evolved from classical matrix factorization techniques to sophisticated tensor algebra frameworks capable of handling arbitrary N-way data structures. Canonical Polyadic Decomposition (CPD), also known as CANDECOMP/PARAFAC, represents the foundational tensor decomposition method. Introduced independently by Carroll and Chang [carroll1970analysis] and Harshman [harshman1970foundations], CPD decomposes a tensor into a sum of rank-one tensors, providing a unique factorization under mild conditions. Tucker decomposition, proposed by Tucker [tucker1966some], provides a more flexible alternative through its core tensor and factor matrices structure. Unlike CPD, Tucker decomposition allows different ranks along each mode, making it suitable for data with varying complexity across dimensions. The Higher-Order SVD (HOSVD) algorithm by De Lathauwer et al. [de2000multilinear] established efficient computational procedures for Tucker decomposition, enabling its application to large-scale problems. Recent advances in tensor decomposition have focused on addressing practical challenges such as missing data, noise robustness, and computational scalability. Tensor completion methods, including work by Liu et al. [liu2012tensor] and Yuan and Zhang [yuan2016tensor], enable handling of incomplete tensors through low-rank assumptions. Robust tensor decomposition techniques [huang2014robust] provide resilience against outliers and corrupted entries. Non-negative tensor factorization (NTF) has gained attention for applications requiring interpretable, parts-based representations. Lee and Seung’s pioneering work [lee2000algorithms] on non-negative matrix factorization was extended to tensors by Shashua and Hazan [shashua2005non], enabling decompositions with non-negativity constraints that often yield more interpretable factors. The integration of tensor decomposition with machine learning has led to tensorized neural networks [novikov2015tensorizing], tensor-based kernel methods [signoretto2011kernel], and multi-task learning frameworks [wimalawarne2014multitask]. These approaches leverage tensor structure to reduce model complexity while maintaining representational power. In the context of clustering, tensor-based methods have shown particular promise for multi-view and multi-way data analysis. Zhou et al. [cichocki2015tensor] developed tensor-based multiway arrays for bid data sets, while Benjamin et al. [benjamin2024tensor] proposed tensorized possibilistic c-means for high-dimensional clustering. However, these methods are not designed for federated environments and lack the privacy-preserving mechanisms required for distributed deployment.
3 Mathematical Preliminaries
This section presents a thorough mathematical formulation of the fundamental concepts and notations employed in this study. The section commences with a review of the principles of federated learning and clustering. This is followed by an examination of multi-view clustering techniques. Subsequently, the section explores heat kernel methods and their applications in quantum-inspired learning. Finally, the discussion will turn to tensor decomposition methods relevant to the proposed framework.
3.1 Federated Learning and Clustering
Federated learning, pioneered by McMahan et al. [mcmahan2017communication], enables collaborative machine learning across distributed clients without centralizing raw data. The seminal FedAvg algorithm established the foundation for aggregating local model updates through weighted averaging. However, traditional federated learning primarily focuses on supervised tasks, with federated clustering receiving comparatively less attention despite its practical importance in distributed data analytics.
Konečnỳ et al. [konevcny2016federated] introduced federated optimization as a method to train a centralized model using decentralized data stored across numerous devices, emphasizing communication efficiency due to limited connectivity and privacy concerns. They proposed modifications to existing algorithms to adapt them to federated optimization settings. Li et al. [li2020acceleration] adopted the foundational ideas of federated optimization and extended them by integrating advanced techniques like gradient compression, acceleration, and variance reduction to address communication bottlenecks and improve the scalability of federated learning systems
Early work on federated clustering by Dennis et al. [dennis2021heterogeneity], Ghosh et al. [ghosh2002consensus], and Makhija et al. [makhija2022federated] demonstrated the feasibility of distributed k-means and hierarchical clustering. These approaches typically involve local clustering followed by global aggregation of cluster centers or assignments. However, these methods assume homogeneous data distributions and fail to address the inherent heterogeneity in federated environments.
Recent advances have introduced personalized federated clustering to address client heterogeneity. Caldarola et al. [caldarola2021cluster] proposed adaptive weight schemes for heterogeneous clustering, while Li et al. [li2022hpfl] developed meta-learning approaches for personalized cluster initialization. These works, however, are limited to single-view data and do not leverage the rich multi-modal information available in modern federated systems.
Privacy-preserving aspects of federated clustering have been addressed through differential privacy mechanisms [kairouz2021advances] and secure aggregation protocols [bonawitz2017practical]. Geyer et al. [geyer2017differentially] demonstrated how differential privacy can be integrated into federated clustering while maintaining acceptable utility. However, these privacy mechanisms often come at the cost of clustering accuracy, particularly in high-dimensional settings.
The communication efficiency challenge in federated clustering has been tackled through gradient compression [konecny2016federated], sparsification techniques [aji2017sparse], and model quantization [reisizadeh2020fedpaq]. While these methods reduce communication overhead, they typically do not exploit the inherent structure of multi-view tensor data for more efficient representation and transmission.
3.2 Multi-View Clustering
@@
Multi-view clustering leverages complementary information from multiple data representations to improve clustering performance. Early work by Bickel and Scheffer [bickel2004multi] established the theoretical foundations for multi-view learning, demonstrating superior performance when views provide complementary information about the underlying data structure.
@@
Classical multi-view clustering approaches can be categorized into three main paradigms: co-training based methods, subspace learning approaches, and consensus clustering techniques. Co-training methods, such as the work by Kumar et al. [kumar2011co], iteratively refine clustering results by leveraging inter-view consistency. Subspace learning approaches, including Canonical Correlation Analysis (CCA) based methods [hotelling1992relations] and shared subspace clustering [wang2019multiple], project multiple views into a common representation space for unified clustering.
@@
Graph-based multi-view clustering has gained significant attention due to its ability to capture complex inter-view relationships. Wang et al. [wang2019gmc] proposed multi-view graph clustering with adaptive neighbor learning, while Nie et al. [nie2016parameter] developed parameter-free auto-weighted multiple graph learning. These methods construct view-specific graphs and learn optimal view weights for clustering.
@@
Deep learning approaches have revolutionized multi-view clustering through end-to-end representation learning. Zhang et al. [zhang2019ae2] introduced auto-encoder based multi-view clustering, while Zhu et al. [zhu2019multi] developed deep multi-view clustering networks with view correlation discovery. However, these methods typically require centralized data access and are not designed for federated environments.
@@
Fuzzy clustering extensions to multi-view scenarios have been explored by Jiang et al. [jiang2014collaborative] and Wang et al. [wang2017multi], introducing weighted aggregation schemes and view-specific fuzzification parameters. These soft clustering approaches provide uncertainty quantification but lack the scalability and privacy considerations required for federated deployment.
@@
The objective is to provide a comprehensive mathematical objective function that can effectively handle multi-view data distributed across multiple clients using tensor algebra, taking into account non-linearity compatibilities. Consequently, the distance between data points and cluster centers across data views in Euclidean space is replaced by a kernel Euclidean distance (KED). The following definition is proposed for the mathematical expression of KED:
@@
As illustrated in Eq. 5, it is imperative to prioritize all data views equally. This approach is not conducive to the efficient resolution of multiple resources data. Given that divergent perspectives correspond to disparate manifestations of feature behavior, it becomes imperative to allocate a distinct weight, denoted as $v_{h}$, to each data view, denoted as $x_{i}^{h}$. It is imperative to underscore that this allocation must be completed, in accordance with the principle that $\sum_{h=1}^{s}v_{h}=1$. By incorporating the weight factors, denoted by $v_{h}$, into the analysis, a modification to the Eq. 5 can be formulated as follows: